OpenStudy (anonymous):

Solve the system of equations by using matrix equations: 5x+4y=-3 -3x-5y=-24

OpenStudy (anonymous):

I forgot all about matrices and how to find the inverse..

hero (hero):

All you joe...

OpenStudy (saifoo.khan):

-_-

hero (hero):

Apparently, nobody likes matrices

hero (hero):

That includes you Saifoo

OpenStudy (anonymous):

Too tedious..

OpenStudy (saifoo.khan):

agree.

OpenStudy (anonymous):

yeah, its just tedious. Use this formula to get the inverse of the matrix: If A is the 2x2 matrix a b c d Then: \[A^{-1}=\frac{1}{detA}B\] Where B is the 2x2 matrix d -b -c a

OpenStudy (anonymous):

det?

OpenStudy (anonymous):

Determinant? if you dont know how to calculate the determinant of a 2x2 matrix then this is going to require a bit more work.

OpenStudy (anonymous):

(not saying that as an insult, im just saying, other paths are long, sry if it came off that way)

OpenStudy (anonymous):

Ah, no it's fine. It's like cross multiplying or something.. right?

OpenStudy (anonymous):

right right, if A is the 2x2 matrix: a b c d then the det(A) = ad - bc

OpenStudy (anonymous):

In this case, the determinant is -37?

OpenStudy (anonymous):

From Mathematica: \[\text{LinearSolve}\left[\left( \begin{array}{cc} 5 & 4 \\ -3 & -5 \end{array} \right),\left( \begin{array}{c} -3 \\ -24 \end{array} \right)\right]\to \left\{\left\{-\frac{111}{13}\right\},\left\{\frac{129}{13}\right\}\right\} \]

OpenStudy (anonymous):

it would be:\[5(-5)-(4)(-3)=-25+12=-13\]

OpenStudy (anonymous):

Oh yeah, my mistake.. I wrote it as -12 instead of +

OpenStudy (anonymous):

\[-1/13\left[\begin{matrix}-5 & -4 \\ 3 & 5\end{matrix}\right]\left[\begin{matrix}5 & 4 \\ 3 & -5\end{matrix}\right] \left(\begin{matrix}x \\ y\end{matrix}\right) = -1/13\left[\begin{matrix}-5 & -4 \\ 3 & 5\end{matrix}\right]\left(\begin{matrix}-3 \\ 24\end{matrix}\right)\] The ( ) are supposed to be [ ] but I couldn't find the [ ] for 2x1 matrices

OpenStudy (anonymous):

@@ I forgot how to multiply matrices too!! How did you get 111? -5(-3).. and then what else?

OpenStudy (anonymous):

its -5(-3)+(-4)(-24) for the first entry, then (3)(-3)+(5)(-24)

OpenStudy (anonymous):

for the second.

OpenStudy (anonymous):

Okay. How did you cancel out everything on the left side so you are just left with x, y ?

OpenStudy (anonymous):

Well i'm not sure exactly which matrix equations you are supposed to use but here's gaussian elimination Original Matrix: 5 4 -3 -3 -5 -24 3 * row 1: 15 12 -9 -3 -5 -24 5 * row 2: 15 12 -9 -15 -25 -120 row 1 added to row 2: 15 12 -9 0 -13 -129 row 2 divided by -13: 15 12 -9 0 1 9.923 -12 * row 2: 15 12 -9 0 -12 -119.076 row 2 added to row 1: 15 0 -128.076 0 -12 -119 row 1 divided by 15 and row 2 divided by -12: 1 0 -8.538 0 1 9.923 therefore: x = -8.538 y = 9.923

OpenStudy (anonymous):

Do you mean the part where I did:\[A\vec x=\vec b\iff A^{-1}A\vec x=A^{-1}\vec b\]\[\vec x = A^{-1}\vec b\] ?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Thats because:\[AA^{-1}=A^{-1}A=I\] The identity matrix. Its the equivalent of 1 in matrices. So guess the step I skipped was: \[A\vec x =\vec b \iff (A^{-1}A)\vec x =A^{-1}\vec b \iff I\vec x = A^{-1}\vec b\iff \vec x = A^{-1}\vec b\]

OpenStudy (anonymous):

Thats the purpose of A inverse. it multiply with A to "cancel it out"

OpenStudy (anonymous):

Oh, I see now! If I could give you more than one medal, I would!!